OpenStudy (anonymous):

Find the product. (u + v)^2 help please!

OpenStudy (zehanz):

(u+v)²=(u+v)(u+v) Now FOIL and add like terms. Better: These kind of things always give the same answer. If you learn them by heart, you can do it much quicker (mostly it is given as (a+b)²= ....)

OpenStudy (anonymous):

u2+ 2uv+v u2+ 2uv+v2 u2+ 2uv+ 2v2 these are my options

OpenStudy (anonymous):

i dont know which one though

OpenStudy (anonymous):

do you know how to FOIL?

OpenStudy (zehanz):

You don't need options. You have to do this yourself. FOIL means: First: multiply first terms: u*u=u² Outer: multiply outer terms: u*v=uv Inner: multiply inner terms: v*u=uv Last: multiply last terms: v*v=v² Now add up: u²+uv+uv+v²=u²+2uv+v²

OpenStudy (anonymous):

yes. but i keep getting something else

OpenStudy (anonymous):

the answer is 1

OpenStudy (anonymous):

@zachy1234 thanks man

OpenStudy (zehanz):

Don't thank him for a wrong answer...

OpenStudy (zehanz):

Did you read my explanation?

OpenStudy (anonymous):

it's not the first one at all .

OpenStudy (zehanz):

@juliajb: That's what I said ;)

OpenStudy (anonymous):

@ZeHanz yes. it doesnt make sence. just math all together. but im doing my assignment. this is my last one.

OpenStudy (zehanz):

It really does make sense, but you have to try and learn from it. You don't get it just by magic...

OpenStudy (anonymous):

@ZeHanz just did it. is it the second one?

OpenStudy (anonymous):

yes! :)

OpenStudy (zehanz):

Now you're talking!

OpenStudy (anonymous):

thanks :)

OpenStudy (zehanz):

yw!